How to solve greatest common divisor ( g.c.d. )of 1125 any 574

Grade:12th pass

4 Answers

Aditya Gupta
4 years ago
We first perform the prime factorisation:
1125= 5^3*3^2
and 574= 2*7*41
Clearly there are no common prime factors. Hence the HCF or greatest common divisor (gcd) is equal to 1.
KINDLY APPROVE :))
Vikas TU
4 years ago
https://www.askiitians.com/revision-notes/class-6-maths/playing-with-numbers/Dear studentTo understand the processPlease find the link below
Good Luck 
Arun
4 years ago
Dear student
 
1125 = 3 * 3 * 5 *5 *5
574 = 2* 7 * 41
 
Hence HCF = G CD = 1 ( as nothing is common)
 
Hope is helpful
Yashraj
3 years ago
i will again say just go through it once again 
the concept is to factorise them and then take common part that is gcd